# Motives of nullcones of quiver representations

### v2

@@ 4.1. Introduction
(2) $\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\tilde{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}~{}.$
Now we extend the index set $\tilde{L}_{\mathbf{d}}$ of dominant and minimal string to the index set $\bar{L}_{\mathbf{d}}$ of dominant and balanced strings. The following theorem shows that this change of index sets does not affect the formula of the motive of the nullcone.
@@
For a balanced string $\mu$ which is not minimal, the semistable locus $R_{d_{s}}^{\theta_{s}-sst}(Q_{s})$ is empty.
@@
For a segment $K=J_{i}$, we define $Q_{K}=Q_{i}$ and ${\bf d}_{K}={\bf d}_{i}$. First note that we can switch from the stability $\theta_{s}$ to the stability $\tilde{\theta}$ given by $\tilde{\theta}(v,j)=-(p_{i}+j)$ on the level quiver $Q_{K}$ for a segment $K$ starting with $p_{i}$, since the open subset of semistable representations does not change when passing from $\theta$ to $x\cdot\theta$ for a positive $x$. We also have $\tilde{\theta}({\bf d}_{K})=0$ for all segments $K$
@@
Moreover, we note that for a fixed string $s=(s_{1},\dots,s_{n})\in\mathbb{Q}^{n}$, the minimum of $\{|s-q|:=\sum_{i=1}^{n}(s_{i}-q)^{2}\}$ for $q\in\mathbb{Q}$ is given by $|s-\bar{q}|$ such that $\sum_{i=1}^{n}(s_{i}-\bar{q})=0$. If, additionally, $\sum_{i=1}^{n}s_{i}<\sum_{i=1}^{n}(s_{i}-q)<0$, then $|s|\geq|s-q|$, since we can consider $|s-q|$ (for a fixed $s$) as a quadratic function in $q$. Analogously, if $\sum_{i=1}^{n}s_{i}>\sum_{i=1}^{n}(s_{i}-q)>0$, then $|s|\geq|s-q|$.
@@
Let $\mu=(\mu_{1},\dots,\mu_{d})$ be a balanced, but not minimal string. By definition, there exists a (balanced) $\mu^{\prime}$ with $|\mu^{\prime}|<|\mu|$ and $R(\mu^{\prime})\supset R(\mu)$. Therefore we find a segment $J=(\mu_{i})_{i\in I}$ in $\mu$ such that the corresponding substring $J^{\prime}=(\mu^{\prime})_{i\in I}$ of $\mu^{\prime}$ satisfies $|J^{\prime}|=\sum_{I}\mu_{i}^{\prime 2}<\sum_{I}\mu_{i}^{2}=|J|$. We can decompose the indexing set $I$ in three disjoint sets given by
$I^{n}_{1}=\{i\in I\text{ with }\mu_{i}>\mu_{i}^{\prime}\}~{},$
$I^{p}_{1}=\{i\in I\text{ with }\mu_{i}<\mu_{i}^{\prime}\}~{},$
$I_{0}=\{i\in I\text{ with }\mu_{i}=\mu_{i}^{\prime}\}~{}.$
@@
Since $|J^{\prime}|<|J|$, either $\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ or $\sum_{I^{p}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{p}_{1}}\mu_{i}^{2}$. Without loss of generality we assume the inequality holds for $I^{n}_{1}$. The other case follows analogously.
@@
Consider $\mu_{i}^{\prime}$ with $i\in I^{n}_{1}$. Then there exists an $\epsilon_{i}>0$ with $\mu_{i}^{\prime}=\mu_{i}-\epsilon_{i}$. In the level quiver $Q_{J}$, there exists a vertex $v_{i}$ corresponding to $\mu_{i}$. Every arrow $\alpha$ in $Q_{J}$ with $t(\alpha)=v_{i}$ corresponds to (possibly several) elements $\mu_{l}\in J$ such that $\mu_{i}-\mu_{l}=1$ and $\pi_{li}\in\Pi$. Since $R(\mu^{\prime})\supset R(\mu)$, every $\mu_{l}$ obtained in the above way fulfills $\mu^{\prime}_{l}=\mu_{l}-\epsilon_{l}$ with $\epsilon_{l}>\epsilon_{i}$. This shows that $I^{n}_{1}$ is left-complete.
For a subset $\hat{I}$ of $I$ we can now construct a subquiver $\hat{Q}_{J}$ of $Q_{J}$ from the substring $\hat{J}$ of $J$ determined by $\hat{I}$ in the same way the level quiver is constructed from a balanced string; we can also construct a dimension vector $\hat{{\bf d}}_{J}$ for this subquiver which fulfills $(\hat{{\bf d}}_{J})_{i}\leq({\bf d}_{J})_{i}$.
We now assume that the semistable locus in question is not empty. We then claim that for all left-complete subsets $\hat{I}$, the inequality $\sum_{\hat{I}}\mu_{i}\leq 0$ holds. Assume, to the contrary, that $\sum_{\hat{I}}\mu_{i}>0$. Let $V$ be an arbitrary representation of dimension vector ${\bf d}_{J}$ of $Q_{J}$. Since $\hat{I}$ is left-complete,we have $({\bf d}_{J}-\hat{\bf d}_{J})_{k}=0$ for all vertices $v_{k}$ in $\hat{Q}_{J}$, which are not a sink in $\hat{Q}_{J}$, and $({\bf d}_{J}-\hat{\bf d}_{J})_{k}=({\bf d}_{J})_{k}$ for all $v_{k}$ which are not in $\hat{Q}_{J}$. Therefore, $V$ admits a subrepresentation $V^{\prime}$ of dimension vector ${\bf d}_{J}-\hat{\bf d}_{J}$. We have $\tilde{\theta}(d_{J})=0$ since $\mu$ is balanced, showing that $\tilde{\theta}(d_{J}-\hat{d}_{J})>0$. This proves the claim.
@@
We know that $\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}=\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ and $\sum_{I^{n}_{1}}\mu_{i}\leq 0$, because $I^{n}_{1}$ is left-complete. We choose $\tilde{\epsilon}_{1}$ as the minimum of all $\epsilon_{i}$ for $i\in I^{n}_{1}$. We get
$\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}<\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}\leq\sum_{I^{n}_{1}}\mu_{i}\leq 0~{},$
leading to $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}$. Now we define
$I^{n}_{2}=\{i\in I^{n}_{1}:\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}<\mu_{i}\}~{}.$
@@
In other words, we exclude all $\mu_{i}^{\prime}$ with $\mu_{i}^{\prime}+\tilde{\epsilon}_{1}=\mu_{i}$. Now $I^{n}_{2}$ is left-complete by the same argument as for $I^{n}_{1}$. We choose $\tilde{\epsilon}_{2}$ as the minimum of all $\epsilon_{i}-\tilde{\epsilon}_{1}$ with $i\in I^{n}_{2}$. We get:
@@
$\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}~{}.$
We get a descending chain of subsets
$I_{1}^{n}\supsetneq I_{2}^{n}\supsetneq\dots~{},$
which has to terminate after finitely many steps. This yields the following estimate: $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>$
$>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{1}\setminus I^{n}_{2}}\mu_{i}^{2}+\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}>\ldots>\sum_{I^{n}_{1}}\mu_{i}^{2}~{}.$
Since we have started with $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$, we have thus constructed a contradiction, proving that a balanced, but not minimal string, has an empty semistable locus.
@@
For the nullcone $N_{\bf d}$ we have the following equation of motives:
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\bar{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}~{}.$
@@
We illustrate these constructions and the calculation of the motive of $N_{\bf d}$ for the $m$-loop quiver for $d=2$ and $d=3$. For $d=2$ the list of strings consists of $(-\frac{1}{2},\frac{1}{2})$ and $(0,0)$. The corresponding quivers $Q_{s}$ are:
@@
We calculate the individual summands in Formula (2) starting with the string $s=(-\frac{1}{2},\frac{1}{2})$ and thus ${\bf d}_{s}=(1,1)$: the space $V_{s}$ consists of $m$-tuples of matrices with one arbitrary entry, consequently $\dim V_{s}=m$. Moreover, the dimension of the representation space $R_{{\bf d}_{s}}(Q_{s})$ equals the number of arrows. The group $P_{s}$ consists of an invertible $2\times 2$-matrix with arbitrary entries at $3$ positions and $\dim G_{d_{s}}=\dim(G_{1}\times G_{1})=2$. We thus find:
$\phi((-\frac{1}{2},\frac{1}{2}))=m-m-3+2=-1~{},$
and similarly
$\phi((0,0))=0-0-4+4=0~{}.$
For the first quiver every representation is semistable except the one consisting only of trivial maps. For the second quiver there is only one representation, which is in addition semistable. Therefore the corresponding motives of semistable loci are $\mathbb{L}^{m}-1$ and $1$, respectively. Consequently, the motive of $L_{2}$ calculates to
$[L_{2}]$ $=(\mathbb{L}^{-1}\frac{(\mathbb{L}^{m}-1)}{[G_{1}]^{2}}+\frac{1}{[G_{2}]})[G_{2}]$
$=(\mathbb{L}^{m-1}-\mathbb{L}^{-1})(\mathbb{L}^{2}+\mathbb{L})+1$
$=\mathbb{L}^{m+1}+\mathbb{L}^{m}-\mathbb{L}~{}.$
The next example is the $m$-loop quiver with dimension vector $d=3$. The relevant strings in the classification are:
$(-1,0,1),(-\frac{2}{3},\frac{1}{3},\frac{1}{3})(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})(-\frac{1}{2},0,\frac{1}{2}),(0,0,0)$
@@
and stability
$(1,0,-1),(2,-1),(1,-2),(1,-1,0),(0)~{}.$
Similarly to the previous example we find
$\phi((-1,0,1))$ $=3m-2m-6+3=m-3~{},$
$\phi((-\frac{2}{3}),\frac{1}{3},\frac{1}{3})$ $=2m-2m-7+5=-2~{},$
$\phi((-\frac{1}{3},-\frac{1}{3},\frac{2}{3},))$ $=2m-2m-7+5=-2~{},$
$\phi((-\frac{1}{2},0,\frac{1}{2}))$ $=m-m-6+3=-3~{},$
$\phi((0,0,0))$ $=0-0-9+9=0~{}.$
It remains to describe the motives of the semistable loci. For the first quiver all representations are semistable unless one tuple of parallel arrows contains only trivial maps. Consequently the resulting motive is $(\mathbb{L}^{m}-1)^{2}$. For the second and third quiver we can write the $m$-tuple of vectors of length $2$ in a $2\times m$-matrix, obtaining that the representation is semistable if and only if this matrix is of rank $2$. Consequently, the motive equals $(\mathbb{L}^{m}-1)(\mathbb{L}^{m}-\mathbb{L})$. The motives for the final two quivers are $\mathbb{L}^{m}-1$ and $1$, respectively, with similar arguments as before.
Putting all this information together, the motive of $N_{3}$ is given by:
$[N_{3}]$ $=(\mathbb{L}^{m-3}\frac{(\mathbb{L}^{m}-1)^{2}}{[G_{1}]^{3}}+2\mathbb{L}^{-2}\frac{(\mathbb{L}^{m}-1)(\mathbb{L}^{m}-\mathbb{L})}{[G_{1}][G_{2}]}+\mathbb{L}^{-3}\frac{\mathbb{L}^{m}-1}{[G_{1}]^{3}}+\frac{1}{[G_{3}]})[G_{3}]$
$=(\mathbb{L}^{3m-3}-2\mathbb{L}^{2m-3}+\mathbb{L}^{m-3})(\mathbb{L}^{6}+2\mathbb{L}^{5}+2\mathbb{L}^{4}+\mathbb{L}^{3})$
$\;\;\;\;+(2\mathbb{L}^{2m-2}-2\mathbb{L}^{m-1}-2\mathbb{L}^{m-2}+2\mathbb{L}^{-1})(\mathbb{L}^{4}+\mathbb{L}^{3}+\mathbb{L}^{2})$
$\;\;\;\;+(\mathbb{L}^{m-3}-\mathbb{L}^{-3})(\mathbb{L}^{6}+2\mathbb{L}^{5}+2\mathbb{L}^{4}+\mathbb{L}^{3})+1$
@@
$=\mathbb{L}^{3m+3}+2\mathbb{L}^{3m+2}+2\mathbb{L}^{3m+1}+\mathbb{L}^{3m}-2\mathbb{L}^{2m+3}-2\mathbb{L}^{2m+2}-2\mathbb{L}^{2m+1}+\mathbb{L}^{3}~{},$
in accordance with example 3.6.
4.2. Rational/integral level quiver
We know that the elements $s$ of the index set ${\bar{L}_{\mathbf{d}}}$ give rise to quivers $Q_{s}$, such that the corresponding Hesselink stratum is non-empty if and only if the semi-stable locus is non-empty. Our next aim is now to define a so-called rational level quiver which combines the level quivers associated to the segments, in such a way that the previous properties are still preserved, but the index set for the Hesselink stratification can be described in terms of specific dimension vectors $\mathbf{e}$ for this fixed quiver.

### v3

@@ 4.1 Introduction
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\widetilde{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{\big[R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})\big]}{[G_{\mathbf{d}_{s}}]}.$ (4.2)
Now we extend the index set $\widetilde{L}_{\mathbf{d}}$ of dominant and minimal strings to the index set $\overline{L}_{\mathbf{d}}$ of dominant and balanced strings. Theorem 4.12 below shows that this change of index sets does not affect the formula for the motive of the nullcone. We need some preparation in advance.
Lemma 4.9.
Let $\mu$ be a balanced string and $J_{i}$ a segment of $\mu$ starting with $p_{i}$. We consider the stability $\tilde{\theta}$ defined by $\tilde{\theta}(v,j)=-(p_{i}+j)$. Then for the dimension vector $\mathbf{d}_{J_{i}}$ we have $\tilde{\theta}(\mathbf{d}_{J_{i}})=0$.
Proof.
We have
$\tilde{\theta}(\mathbf{d}_{J_{i}})=\sum_{(v,j)\in Q_{J_{i}}}-a_{i,k}^{v}(p_{j}+j)=\sum_{j=1}^{k_{i}}a_{i,k}(p_{j}+j)=\sum_{k\in\tilde{J}_{i}}\mu_{k}=0,$
because $\mu$ is balanced. ∎
Definition 4.10.
For a segment $K$, we define $Q_{K}=Q_{i}$ and ${\bf d}_{K}={\bf d}_{i}$ if $K=J_{i}$.
Let $\mu$ be a string and $J$ a segment of $\mu$ with corresponding index set $I$. For a subset $\hat{I}$ of $I$, we can construct a subquiver $\hat{Q}_{J}$ of $Q_{J}$ from the substring $\hat{J}$ of $J$ determined by $\hat{I}$ in the same way the level quiver is constructed from a balanced string; we can also construct a dimension vector $\hat{{\bf d}}_{J}$ for this subquiver which fulfills $\bigl(\hat{{\bf d}}_{J}\bigr)_{i}\leq({\bf d}_{J})_{i}$.
Lemma 4.11.
If the semistable locus $R_{d_{J}}^{\tilde{\theta}\text{\rm-sst}}(Q_{J})$ is non-empty, then for all left-complete subsets $\hat{I}$ of the index set $I$ corresponding to the segment $J$ the inequality $\sum_{\hat{I}}\mu_{i}\leq 0$ holds.
Proof.
Assume, to the contrary, that $\sum_{\hat{I}}\mu_{i}>0$. Let $V$ be an arbitrary representation of dimension vector ${\bf d}_{J}$ of $Q_{J}$. Since $\hat{I}$ is left-complete,we have $\bigl({\bf d}_{J}-\hat{\bf d}_{J}\bigr)_{k}=0$ for all vertices $v_{k}$ in $\hat{Q}_{J}$, which are not a sink in $\hat{Q}_{J}$, and $\bigl({\bf d}_{J}-\hat{\bf d}_{J}\bigr)_{k}=({\bf d}_{J})_{k}$ for all $v_{k}$ which are not in $\hat{Q}_{J}$. Therefore, $V$ admits a subrepresentation $V^{\prime}$ of dimension vector ${\bf d}_{J}-\hat{\bf d}_{J}$. By Lemma 4.9, we have $\tilde{\theta}(\mathbf{d}_{J})=0$ since $\mu$ is balanced, showing that $\tilde{\theta}\bigl(\mathbf{d}_{J}-\hat{\bf d}_{J}\bigr)>0$. This proves the claim. ∎
@@
For a balanced string $s$ which is not minimal, the semistable locus $R_{{\bf d}_{s}}^{\theta_{s}\text{\rm-sst}}(Q_{s})$ is empty.
@@
First note that we can switch from the stability $\theta_{s}$ to the stability $\tilde{\theta}$ given by $\tilde{\theta}(v,j)=-(p_{i}+j)$ on the level quiver $Q_{K}$ for a segment $K$ starting with $p_{i}$, since the open subset of semistable representations does not change when passing from $\theta$ to $x\cdot\theta$ for a positive $x$.
@@
Moreover, we note that for a fixed string $s=(s_{1},\dots,s_{n})\in\mathbb{Q}^{n}$, the minimum of $\big\{|s-q|:=\sum_{i=1}^{n}(s_{i}-q)^{2}\big\}$ for $q\in\mathbb{Q}$ is given by $|s-\bar{q}|$ such that $\sum_{i=1}^{n}(s_{i}-\bar{q})=0$. If, additionally, $\sum_{i=1}^{n}s_{i}<\sum_{i=1}^{n}(s_{i}-q)<0$, then $|s|\geq|s-q|$, since we can consider |sq||s-q| (for a fixed $s$) as a quadratic function in $q$. Analogously, if $\sum_{i=1}^{n}s_{i}>\sum_{i=1}^{n}(s_{i}-q)>0$, then $|s|\geq|s-q|$.
@@
Let $\mu=(\mu_{1},\dots,\mu_{d})$ be a balanced, but not minimal string. By definition, there exists a (balanced) $\mu^{\prime}$ with $|\mu^{\prime}|<|\mu|$ and $\mathcal{R}(\mu^{\prime})\supset\mathcal{R}(\mu)$. Therefore, we find a segment $J=(\mu_{i})_{i\in I}$ in $\mu$ such that the corresponding substring $J^{\prime}=(\mu^{\prime})_{i\in I}$ of $\mu^{\prime}$ satisfies $|J^{\prime}|=\sum_{I}\mu_{i}^{\prime 2}<\sum_{I}\mu_{i}^{2}=|J|$. We can decompose the indexing set $I$ in three disjoint sets given by
$I^{n}_{1}=\{i\in I\text{ with }\mu_{i}>\mu_{i}^{\prime}\},\qquad I^{p}_{1}=\{i\in I\text{ with }\mu_{i}<\mu_{i}^{\prime}\},\qquad I_{0}=\{i\in I\text{ with }\mu_{i}=\mu_{i}^{\prime}\}.$
@@
Since $|J^{\prime}|<|J|$, either $\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ or $\sum_{I^{p}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{p}_{1}}\mu_{i}^{2}$. Without loss of generality, we assume the inequality holds for $I^{n}_{1}$. The other case follows analogously.
@@
Consider $\mu_{i}^{\prime}$ with $i\in I^{n}_{1}$. Then there exists an $\epsilon_{i}>0$ with $\mu_{i}^{\prime}=\mu_{i}-\epsilon_{i}$. In the level quiver $Q_{J}$, there exists a vertex $v_{i}$ corresponding to $\mu_{i}$. Every arrow $\alpha$ in $Q_{J}$ with $t(\alpha)=v_{i}$ corresponds to (possibly several) elements $\mu_{l}\in J$ such that $\mu_{i}-\mu_{l}=1$ and $\pi_{li}\in\Pi$. Since $\mathcal{R}(\mu^{\prime})\supset\mathcal{R}(\mu)$, every $\mu_{l}$ obtained in the above way fulfills $\mu^{\prime}_{l}=\mu_{l}-\epsilon_{l}$ with $\epsilon_{l}>\epsilon_{i}$. This shows that $I^{n}_{1}$ is left-complete.
@@
We now assume that the semistable locus in question is not empty. By Lemma 4.11, we know that for all left-complete subset $\hat{I}$ the inequality $\sum_{\hat{I}}\mu_{i}\leq 0$ hold. We know that $\smash{\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}}=\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ and $\sum_{I^{n}_{1}}\mu_{i}\leq 0$, because $I^{n}_{1}$ is left-complete. We choose $\tilde{\epsilon}_{1}$ as the minimum of all $\epsilon_{i}$ for $i\in I^{n}_{1}$. We get
$\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}<\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}\leq\sum_{I^{n}_{1}}\mu_{i}\leq 0,$
@@
leading to $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}$. Now we define $I^{n}_{2}=\{i\in I^{n}_{1}\mid\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}<\mu_{i}\}$. In other words, we exclude all $\mu_{i}^{\prime}$ with $\mu_{i}^{\prime}+\tilde{\epsilon}_{1}=\mu_{i}$. Now $I^{n}_{2}$ is left-complete by the same argument as for $I^{n}_{1}$. We choose $\tilde{\epsilon}_{2}$ as the minimum of all $\epsilon_{i}-\tilde{\epsilon}_{1}$ with $i\in I^{n}_{2}$. We get
@@
$\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}.$
We get a descending chain of subsets $I_{1}^{n}\supsetneq I_{2}^{n}\supsetneq\cdots$, which has to terminate after finitely many steps. This yields the following estimate:
$\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{1}\setminus I^{n}_{2}}\mu_{i}^{2}+\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}>\dots>\sum_{I^{n}_{1}}\mu_{i}^{2}.$
Since we have started with $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$, we have thus constructed a contradiction, proving that a balanced, but not minimal string, has an empty semistable locus. ∎
@@
For the nullcone $N_{\bf d}$, we have the following equality of motives:
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\overline{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{\big[R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})\big]}{[G_{\mathbf{d}_{s}}]}.$
@@
We illustrate these constructions and the calculation of the motive of $N_{d}$ for the $m$-loop quiver for $d=2$ and $d=3$. For $d=2$, the list of strings consists of $\bigl(-\frac{1}{2},\frac{1}{2}\bigr)$ and $(0,0)$. The corresponding quivers $Q_{s}$ are
@@
We calculate the individual summands in formula (4.2) starting with the string $s=\bigl(-\frac{1}{2},\frac{1}{2}\bigr)$ and thus ${\bf d}_{s}=(1,1)$: the space $V_{s}$ consists of $m$-tuples of matrices with one arbitrary entry, consequently $\dim V_{s}=m$. Moreover, the dimension of the representation space $R_{{\bf d}_{s}}(Q_{s})$ equals the number of arrows. The group $P_{s}$ consists of an invertible $2\times 2$ matrix with arbitrary entries at $3$ positions and $\dim G_{d_{s}}=\dim(G_{1}\times G_{1})=2$. We thus find
$\phi\left(\left(-\frac{1}{2},\frac{1}{2}\right)\right)=m-m-3+2=-1,$
and similarly $\phi((0,0))=0-0-4+4=0$. For the first quiver, every representation is semistable except the one consisting only of trivial maps. For the second quiver, there is only one representation, which is in addition semistable. Therefore, the corresponding motives of semistable loci are $\mathbb{L}^{m}-1$ and $1$, respectively. Consequently, the motive of $N_{2}$ calculates to
$[N_{2}]$ $=\left(\mathbb{L}^{-1}\frac{(\mathbb{L}^{m}-1)}{[G_{1}]^{2}}+\frac{1}{[G_{2}]}\right)[G_{2}]=\bigl(\mathbb{L}^{m-1}-\mathbb{L}^{-1}\bigr)\bigl(\mathbb{L}^{2}+\mathbb{L}\bigr)+1=\mathbb{L}^{m+1}+\mathbb{L}^{m}-\mathbb{L}.$
The next example is the $m$-loop quiver with dimension vector $d=3$. The relevant strings in the classification are
$(-1,0,1),\left(-\frac{2}{3},\frac{1}{3},\frac{1}{3}\right)\left(-\frac{1}{3},-\frac{1}{3},\frac{2}{3}\right)\left(-\frac{1}{2},0,\frac{1}{2}\right),(0,0,0)$
@@
and stability (1,0,1)(1,0,-1), (2,1)(2,-1), (1,2)(1,-2), (1,1,0)(1,-1,0), $(0)$. Similarly to the previous example, we find
$\phi((-1,0,1))=3m-2m-6+3=m-3,$
$\phi\left(\left(-\frac{2}{3}\right),\frac{1}{3},\frac{1}{3}\right)=2m-2m-7+5=-2,$
$\phi\left(\left(-\frac{1}{3},-\frac{1}{3},\frac{2}{3},\right)\right)=2m-2m-7+5=-2,$
$\phi\left(\left(-\frac{1}{2},0,\frac{1}{2}\right)\right)=m-m-6+3=-3,\qquad\phi((0,0,0))=0-0-9+9=0.$
It remains to describe the motives of the semistable loci. For the first quiver, all representations are semistable unless one tuple of parallel arrows contains only trivial maps. Consequently, the resulting motive is $(\mathbb{L}^{m}-1)^{2}$. For the second and third quiver, we can write the $m$-tuple of vectors of length $2$ in a $2\times m$ matrix, obtaining that the representation is semistable if and only if this matrix is of rank $2$. Consequently, the motive equals $(\mathbb{L}^{m}-1)(\mathbb{L}^{m}-\mathbb{L})$. The motives for the final two quivers are $\mathbb{L}^{m}-1$ and $1$, respectively, with similar arguments as before.
Putting all this information together, the motive of $N_{3}$ is given by
$[N_{3}]={}$ $\left(\mathbb{L}^{m-3}\frac{(\mathbb{L}^{m}-1)^{2}}{[G_{1}]^{3}}+2\mathbb{L}^{-2}\frac{(\mathbb{L}^{m}-1)(\mathbb{L}^{m}-\mathbb{L})}{[G_{1}][G_{2}]}+\mathbb{L}^{-3}\frac{\mathbb{L}^{m}-1}{[G_{1}]^{3}}+\frac{1}{[G_{3}]}\right)[G_{3}]$
$={}$ $\bigl(\mathbb{L}^{3m-3}-2\mathbb{L}^{2m-3}+\mathbb{L}^{m-3}\bigr)\bigl(\mathbb{L}^{6}+2\mathbb{L}^{5}+2\mathbb{L}^{4}+\mathbb{L}^{3}\bigr)$
$+\bigl(2\mathbb{L}^{2m-2}-2\mathbb{L}^{m-1}-2\mathbb{L}^{m-2}+2\mathbb{L}^{-1}\bigr)\bigl(\mathbb{L}^{4}+\mathbb{L}^{3}+\mathbb{L}^{2}\bigr)$
$+\bigl(\mathbb{L}^{m-3}-\mathbb{L}^{-3}\bigr)\bigl(\mathbb{L}^{6}+2\mathbb{L}^{5}+2\mathbb{L}^{4}+\mathbb{L}^{3}\bigr)+1$
@@
$={}$ $\mathbb{L}^{3m+3}+2\mathbb{L}^{3m+2}+2\mathbb{L}^{3m+1}+\mathbb{L}^{3m}-2\mathbb{L}^{2m+3}-2\mathbb{L}^{2m+2}-2\mathbb{L}^{2m+1}+\mathbb{L}^{3},$
in accordance with the example in Section 3.2.
4.2 Rational/integral level quiver
We know that the elements $s$ of the index set $\overline{L}_{\mathbf{d}}$} give rise to quivers $Q_{s}$, such that the corresponding Hesselink stratum is non-empty if and only if the semi-stable locus is non-empty. Our next aim is now to define a so-called rational level quiver which combines the level quivers associated to the segments, in such a way that the previous properties are still preserved, but the index set for the Hesselink stratification can be described in terms of specific dimension vectors $\mathbf{e}$ for this fixed quiver.